The perimeter of a rectangle equals one and a half times its area. Express the length of the rectangle in terms of the width (use variable "w" for width).
Found 2 solutions by nerdybill, solver91311:
Answer by nerdybill(6959)

The perimeter of a rectangle equals one and a half times its area. Express the length of the rectangle in terms of the width (use variable "w" for width)
.
Area of any rectangle = width * length
Perimeter of any rectangle = 2(width+length)
.
Let w = width
and temporarily
x = length
.
From the problem, perimeter is:
1.5wx
.
But,
Perimeter = 2(w+x)
So, then we have
1.5wx = 2(w+x)
Solving for x:
1.5wx = 2(w+x)
1.5wx = 2w+2x
1.5wx-2x = 2w
factoring the left:
x(1.5w-2) = 2w
x = 2w/(1.5w-2)
.
So, your answer would be
length is equal to:
2w/(1.5w-2)

Answer by solver91311(16885)


The area of a rectangle is given by:



and the Perimeter is given by:



Where represents the measure of the length and represents the measure of the width.

But we are given that



So:












John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism